# OPENXRD: A Comprehensive Benchmark and Enhancement Framework for LLM/MLLM XRD Question Answering

### v1

@@ 3.4 Open-Book Mode Observations Using AI-Generated Materials with Expert Review
GPT-4.5-preview -2.77 -3.69 -0.92
O3-mini +0.46 +0.96 +0.50
O1 +0.46 +1.34 +0.88
GPT-4-turbo-preview -0.92 +1.39 +2.31
GPT-4 +0.92 +1.39 +0.47
LLaVA-v1.6-34B +4.61 +11.50 +6.89
LLaVA-onevision-QWEN2-7B-ov-chat +6.45 +10.10 +3.65
LLaVA-onevision-QWEN2-7B-si +5.53 +8.70 +3.17
LLaVA-onevision-QWEN2-7B-ov +5.99 +8.80 +2.81
LLaVA-v1.6-mistral-7B +5.75 +11.10 +5.35
LLaVA-onevision-QWEN2-0.5B-ov +3.68 +6.90 +3.22
LLaVA-v1.5-13B -2.30 +1.90 +4.20
LLaVA-onevision-QWEN2-0.5B-si +5.07 +7.80 +2.73
@@
Large ModelsMid-sized ModelsSmall Models0.370.37-0.37- 0.37%$5.71$%$3.11$%$0.08$%$9.64$%$6.55$%AI-Generated MaterialsExpert-Reviewed Materials
Fig.  7: Average performance gains from AI-generated vs. expert-reviewed supporting materials across model size categories. Large Models (e.g., GPT-4, GPT-4.5, O1, O3-mini); Mid-sized Models (e.g., LLaVA-34B, QWEN2-7B); Small Models (e.g., LLaVA-13B, LLaVA-7B, QWEN2-0.5B). Note the consistently larger gains achieved with expert-reviewed supporting textual materials, particularly evident in mid-sized models.
@@ 4 Discussion
Our research identifies two primary methods for enhancing model performance with domain expertise: directly embedding knowledge into large models, such as GPT-4.5 and O3-mini, and providing models with access to expert-reviewed references during inference. This latter strategy proves especially useful for mid-capacity models, enabling them to nearly match the performance of their larger counterparts by leveraging a "knowledge bridge."
For example, a medium-sized model like LLaVA-v1.6-34B, when paired with expert materials, delivers performance close to that of GPT-4 but with far fewer resources required. The integration of expert-reviewed domain knowledge is not only beneficial in textual analysis but also in multi-modal contexts such as medical imaging or geospatial analysis, where expert-reviewed context can similarly boost specialized question-answering performance and unlock new application areas.
However, the addition of external information is not always beneficial. For instance, GPT-4.5-preview shows a decline in performance when supplemented with external materials, possibly due to conflicts with its pre-trained internal knowledge. This suggests that more information is not necessarily better, particularly for top-tier models that have been extensively trained.
Future research should focus on automated methods to determine when expert review is most beneficial, multi-stage approaches for targeted expert review, and the potential for fine-tuning models to better utilize external references. Extending these methodologies to other specialized domains could further validate the effectiveness of this approach.
@@ 5 Conclusions
We present OPENXRD, an open-book question-answering framework for crystallography that shows how AI-generated supporting materials can improve model performance on specialized scientific questions. Through evaluation on 217 expert-level XRD questions, we find that mid-capacity models (7B-34B parameters) benefit most from external knowledge, gaining up to +11.5 percentage points with expert-reviewed materials compared to +6% with AI-generated content alone. Our main finding is an inverted-U relationship: the largest models show minimal improvement or slight degradation with external materials, while the smallest models have limited capacity to utilize additional information effectively. Mid-sized models occupy a "sweet spot" where they have sufficient processing capability but incomplete domain knowledge. The practical value lies in demonstrating that carefully curated supporting materials can help smaller models approach larger model performance on specialized tasks. This suggests cost-effective deployment strategies where mid-sized models with expert-reviewed materials can substitute for much larger, more expensive models in domain-specific applications. We release our dataset and code to enable further research on how different AI systems utilize external knowledge in specialized domains. While our work focuses on crystallography, the methodology could extend to other technical fields requiring expert-level question-answering.

### v2

@@ 3.5.5 Interpretation: Quality, Not Quantity, Drives Improvements
openai/gpt-5 -1.84 -3.68 -1.84
x-ai/grok-4-fast -0.92 -5.78 -4.86
openai/gpt-5-codex -1.84 -4.14 -2.30
google/gemini-2.5-pro -6.91 -2.76 +4.15
o3-mini -4.15 -3.69 +0.46
meituan/longcat-flash-chat -2.77 -3.23 -0.46
gpt-4.5-preview -2.31 -3.23 -0.92
anthropic/claude-3.5-sonnet -1.84 -1.84 +0.00
openai/gpt-4o -3.18 -3.64 -0.46
dziner-qwen-2.5-72b -3.22 -4.14 -0.92
qwen/qwen3-next-80b-a3b-instruct -0.92 -0.92 +0.00
perplexity/sonar-pro +0.00 -2.30 -2.30
deepseek/deepseek-v3.1-terminus +0.92 +0.92 +0.00
qwen/qwen-plus +0.92 -0.46 -1.38
qwen/qwen3-next-80b-a3b-thinking -2.76 -2.76 +0.00
deepseek/deepseek-chat +1.84 +0.92 -0.92
qwen/qwen-max -1.38 +0.00 +1.38
o1 -1.38 -0.46 +0.92
anthropic/claude-3-opus -0.92 +0.00 +0.92
openai/o1-mini +0.46 +0.00 -0.46
meta-llama/llama-3.1-405b-instruct -4.61 -3.23 +1.38
qwen/qwen-2.5-72b-instruct -1.38 -1.85 -0.47
mistralai/mistral-large -2.31 +0.00 +2.31
amazon/nova-pro-v1 +0.53 +0.53 +0.00
gpt-4-turbo +0.46 +0.46 +0.00
meta-llama/llama-3-70b-instruct -4.14 -1.84 +2.30
meta-llama/llama-3.1-70b-instruct -1.38 +0.40 +1.78
dziner-qwen-2.5-coder-32b +0.92 -6.45 -7.37
openai/gpt-4-0314 -5.99 -0.92 +5.07
gpt-4-turbo-preview -0.92 +1.38 +2.30
amazon/nova-lite-v1 +2.76 +3.68 +0.92
dziner-qwen-2.5-7b -2.31 -2.31 +0.00
gpt-4 +0.46 +0.92 +0.46
openai/gpt-4o-mini +0.92 +0.00 -0.92
openrouter/auto +0.46 +0.46 +0.00
qwen/qwen-2.5-7b-instruct -4.60 -1.38 +3.22
google/gemma-2-27b-it -0.46 +0.46 +0.92
anthropic/claude-3.5-haiku +10.45 +8.35 -2.10
mistralai/mixtral-8x22b-instruct +3.28 +7.08 +3.80
mistralai/mistral-7b-instruct +2.07 +5.30 +3.23
google/gemma-2-9b-it +2.30 +6.45 +4.15
amazon/nova-micro-v1 +4.61 +6.00 +1.39
mistralai/mistral-small +4.61 +8.76 +4.15
anthropic/claude-3-haiku +8.45 +9.29 +0.84
openai/gpt-3.5-turbo-16k -1.59 +3.25 +4.84
mistralai/mixtral-8x7b-instruct +3.63 +4.81 +1.18
meta-llama/llama-3-8b-instruct +1.38 +3.69 +2.31
mistralai/pixtral-12b +3.75 +6.05 +2.30
openai/gpt-3.5-turbo +0.83 +5.66 +4.83
meta-llama/llama-3.1-8b-instruct +6.00 +9.22 +3.22
llava-v1.6-34b +5.99 +11.52 +5.53
lmms-lab/llava-onevision-qwen2-7b-si +5.53 +8.76 +3.23
lmms-lab/llava-onevision-qwen2-7b-ov-chat +6.45 +10.14 +3.69
lmms-lab/llava-onevision-qwen2-7b-ov +5.99 +8.75 +2.76
meta-llama/llama-3.2-3b-instruct -2.31 +0.46 +2.77
arcee-ai/afm-4.5b +2.31 +5.07 +2.76
mistralai/mistral-7b-instruct-v0.1 -5.47 +2.94 +8.41
perplexity/sonar +9.60 +7.90 -1.70
undi95/remm-slerp-l2-13b +4.42 +8.04 +3.62
llamat-3-chat -26.73 -25.34 +1.39
alibaba/tongyi-deepresearch-30b-a3b +24.42 +30.88 +6.46
gryphe/mythomax-l2-13b +7.91 +12.74 +4.83
llava-v1.6-mistral-7b +5.54 +11.07 +5.53
llamat-2-chat -29.56 -34.56 -5.00
lmms-lab/llava-onevision-qwen2-0.5b-ov +3.68 +6.91 +3.23
llava-v1.5-13b -2.30 +1.85 +4.15
lmms-lab/llava-onevision-qwen2-0.5b-si +5.07 +7.83 +2.76
qwen/qwen3-coder-flash +40.87 +42.25 +1.38
qwen/qwen3-coder-plus +65.95 +65.95 +0.00
honeybee-13b +0.92 -0.92 -1.84
qwen/qwen-2.5-coder-32b-instruct +5.98 +9.74 +3.76
honeybee-7b +3.69 -1.84 -5.53
@@
llamat-2 +3.23 +3.23 +0.00
Large ModelsMid-sized ModelsSmall Models1.28-1.28%2.712.71%6.186.18%1.77-1.77%4.444.44%8.528.52%AI-Generated MaterialsExpert-Reviewed Materials
Fig. 10: Average performance gains from AI-generated vs. expert-reviewed supporting materials across model size categories. Large Models (>>70B parameters or advanced architectures like GPT-4, GPT-5, O1, O3-mini, n=27); Mid-sized Models (>>7B-70B parameters including LLaVA-34B, QWEN2-7B, Mistral variants, n=27); Small Models ($\leq$7B parameters including QWEN2-0.5B, LLaMA-3.2-3B, LLaVA-7B, LLaMAT variants, Honeybee-7B, n=20). Note that large models show performance degradation with supporting materials, while mid-sized and small models benefit substantially from expert-reviewed materials. Overall, these results show that as models become larger and contain more built-in knowledge, the additional benefit from external information gradually diminishes.
@@ 3.6 Detailed Subtask-Level Performance with Expert-Reviewed Materials
Mathematically intensive subtasks including Bragg’s Law derivations, structure factor calculations, and reflection condition analysis showed universal 0% improvement with supporting materials across all 74 models. For instance, when asked about structure factors for base-centered lattices with mixed h,k indices, models consistently failed to execute the symbolic computation $F=f[1+e^{i\pi(h+k)}]$ despite correct textual explanations, revealing that current LLMs cannot perform the formal algebraic operations required for crystallographic problem-solving.
@@
3.7 Impact of Token Budget Constraints on Open-Book Performance
While the previous analyses examined model performance under standard inference conditions, we now investigate how computational constraints—specifically, output token budget limitations—affect models’ ability to leverage external knowledge. Token budget restrictions are particularly relevant in deployment scenarios where computational resources or API costs impose practical limits on generation length. Understanding how these constraints interact with open-book augmentation is crucial for real-world applications.
We systematically evaluated two models across varying maximum token budgets (256 to 4096 tokens): llamat-3-chat, a 7B-parameter specialized model, and O3-mini, a reasoning-optimized model designed for scientific tasks. Figure 11 presents the accuracy trends as token budgets increase.
2562565125127687681,0241{,}0241,5361{,}5362,0482{,}0482,5602{,}5603,0723{,}0723,5843{,}5844,0964{,}0962020404060608080100100Max TokensAccuracy (%)(a) llamat-3-chatClosed-bookOpen-book2562565125127687681,0241{,}0241,5361{,}5362,0482{,}0482,5602{,}5603,0723{,}0723,5843{,}5844,0964{,}096707080809090100100Max TokensAccuracy (%)(b) O3-miniClosed-bookOpen-book
Fig. 11: Impact of token budget constraints on closed-book and open-book performance for (a) llamat-3-chat and (b) O3-mini. The x-axis shows maximum output tokens allowed, while the y-axis shows accuracy on the 217-question benchmark. Note the different y-axis scales: llamat-3-chat ranges 20-100% while O3-mini ranges 70-100% due to their different baseline capabilities.
The results reveal markedly different patterns between the two models. For llamat-3-chat (Figure 11a), closed-book performance remains relatively stable across token budgets (55-58%), showing minimal sensitivity to generation length constraints. However, open-book performance exhibits a counterintuitive decline: accuracy drops from 36.87% at 256 tokens to 26.73% at 2048+ tokens—a degradation of approximately 10 percentage points. This inverse relationship suggests severe token starvation: as the model attempts to process longer supporting materials, it exhausts its token budget on input comprehension and internal reasoning, leaving insufficient capacity for coherent answer generation. The supporting materials, rather than aiding reasoning, become a computational burden that crowds out the model’s ability to formulate responses.
In contrast, O3-mini (Figure 11b) demonstrates more favorable scaling properties. Closed-book accuracy improves substantially from 77.42% at 256 tokens to a peak of 95.39% at 2560 tokens, indicating the model benefits from additional reasoning space for complex crystallographic problems. Open-book performance similarly improves from 80.65% to approximately 90%, stabilizing around 1024 tokens. The gap between closed-book and open-book performance narrows at higher token budgets, suggesting that O3-mini can more effectively integrate external knowledge when given adequate computational headroom. However, even O3-mini shows slight performance degradation beyond 2560 tokens in closed-book mode, hinting at potential overfitting or unnecessary elaboration when unconstrained.
These findings have important implications for deployment strategies. First, token budget selection requires careful calibration: smaller models like llamat-3-chat exhibit pathological behavior under open-book conditions when token budgets exceed their processing capacity, while more capable models like O3-mini require minimum token budgets ( 1024 tokens) to effectively leverage external materials. Second, the phenomenon of "open-book degradation" in resource-constrained models suggests that naïvely providing supporting materials without ensuring sufficient token allocation for answer synthesis can be counterproductive. Third, the results underscore fundamental architectural differences: reasoning-optimized models (O3-mini) appear better equipped to manage the token allocation trade-off between comprehending external context and generating accurate responses.
From a practical standpoint, these observations suggest that effective open-book augmentation requires not only high-quality supporting materials but also appropriate computational provisioning. For deployment scenarios with strict token limits, practitioners should either: (1) use models specifically designed for reasoning tasks that can efficiently allocate tokens between context processing and response generation, (2) employ adaptive token budgets that scale with input complexity, or (3) pre-process supporting materials to reduce their token footprint while preserving essential information. Future work should investigate token allocation strategies that dynamically balance between context comprehension and response generation, potentially through attention mechanism modifications or explicit token budgeting protocols.
@@ 4 Discussion
While OPENXRD focuses on inference-time context assimilation, it is explicitly designed to complement parameter-efficient fine-tuning (PEFT) approaches. PEFT methods can encode crystallographic knowledge directly into model weights through techniques like LoRA 39, whereas OPENXRD diagnoses whether models—adapted or not—can effectively utilize answer-guiding context that does not directly reveal solutions. A comprehensive evaluation strategy might compare: (i) closed-book performance after PEFT on external XRD corpora, (ii) open-book performance after PEFT using our oracle passages, and (iii) our baseline open-book performance without PEFT. This three-way comparison decomposes gains attributable to parametric adaptation versus gains from inference-time guidance, a separation difficult to obtain in end-to-end training studies. For instance, a PEFT-adapted model that shows minimal open-book improvement might indicate successful knowledge internalization, while a model showing large open-book gains suggests incomplete adaptation where external guidance remains valuable. OPENXRD standardizes this diagnostic capability across architectures and scales.
@@
Text-based supporting materials prove insufficient for mathematically intensive subtasks requiring symbolic manipulation. Consider a diagnostic case from our benchmark: “For a base-centered unit cell, what is the structure factor F when h and k are mixed (one even, one odd)?” Despite expert-reviewed materials correctly explaining that atoms at (0,0,0) and (1/2,1/2,0) produce destructive interference when h and k have mixed parity (yielding F=0), multiple frontier models incorrectly answered F=2f. Analysis reveals they cannot maintain intermediate symbolic states during phase calculations $\phi=2\pi(hx+ky+lz)$ or correctly apply $F=f[1+e^{i\pi(h+k)}]=0$ when $(h+k)$ is odd. This failure extends to Bragg’s Law derivations, structure factor algebra, and powder diffraction indexing—all showing 0% improvement across our 74-model evaluation regardless of size or architecture.
To overcome this architectural limitation, future work should integrate LLMs with symbolic computation engines such as SymPy 58 or Wolfram Alpha 102 for exact algebraic manipulations, crystallographic knowledge graphs encoding systematic absence rules and space group constraints as structured logical representations rather than text, and domain-specific software including GSAS-II 96 for powder diffraction refinement and Mercury 55 for structure visualization. Such hybrid architectures would enable the LLM to parse problems linguistically, delegate mathematical operations to specialized modules, and synthesize results—combining linguistic flexibility with computational rigor. This approach extends beyond crystallography to other physics-heavy domains requiring both conceptual reasoning and mathematical precision.
@@
Our research identifies two primary methods for enhancing model performance with domain expertise: directly embedding knowledge into large models, such as frontier models like GPT-5, GPT-4.5, and O3-mini, and providing models with access to expert-reviewed references during inference. This latter strategy proves especially useful for mid-capacity models, enabling them to nearly match the performance of their larger counterparts by leveraging a “knowledge bridge.”
For example, a medium-sized model like LLaVA-v1.6-34B, when paired with expert materials, achieves 78.34% accuracy, approaching the performance of much larger models like Llama-3.1-405B (84.33%) and even surpassing some frontier models in cost-effectiveness but with far fewer resources required. The integration of expert-reviewed domain knowledge is not only beneficial in textual analysis but also in multi-modal contexts such as medical imaging or geospatial analysis, where expert-reviewed context can similarly boost specialized question-answering performance and unlock new application areas.
However, the addition of external information is not always beneficial. For instance, multiple frontier models show performance degradation with expert-reviewed supporting materials: GPT-4.5-preview (-3.23%), GPT-5 (-3.68%), Grok-4-fast (-5.78%), Claude-3.5-Sonnet (-1.84%), and dziner-qwen-2.5-72b (-4.14%), possibly due to conflicts with their pre-trained internal knowledge. This suggests that more information is not necessarily better, particularly for top-tier models that have been extensively trained, as the interference from external context can disrupt their already-comprehensive internal reasoning processes.
Our results show that both small and mid-sized models benefit from external knowledge, with small models exhibiting the largest relative gains, whereas large models experience interference. This reveals a fundamental insight: external knowledge augmentation is most effective within a specific capability range where models have sufficient processing power but incomplete domain coverage. This finding has important implications for cost-effective deployment strategies in specialized scientific domains.
Future research should focus on automated methods to determine when expert review is most beneficial, multi-stage approaches for targeted expert review, and the potential for fine-tuning models to better utilize external references, as well as mechanisms to detect when models have sufficient internal knowledge that external context becomes counterproductive.
Beyond incremental improvements, addressing the universal mathematical reasoning failures documented in our benchmark requires architectural innovation. Hybrid systems coupling LLMs with symbolic algebra engines, structured crystallographic knowledge graphs encoding lattice symmetries and systematic absence rules, and specialized software for diffraction simulation and structure refinement would enable models to delegate exact computations while maintaining linguistic problem-solving capabilities. Multi-modal extensions incorporating actual XRD patterns and crystal structure visualizations would test whether vision-language models can integrate visual and textual information. Dynamic retrieval replacing oracle passages with corpus-retrieved content would decompose RAG accuracy into retrieval versus assimilation components. Fine-tuning studies using OPENXRD before and after domain adaptation would reveal whether parametric knowledge embedding complements or supersedes inference-time guidance. Cross-domain validation in related scattering techniques and computational chemistry would test whether our findings about model capacity and external knowledge effectiveness generalize beyond crystallography.
In parallel, we are developing an open-ended XRD evaluation set (short-answer and problem-solving tasks) to complement OPENXRD’s multiple-choice diagnostic benchmark and to enable qualitative assessment of explanation quality and reasoning depth.
@@ 5 Conclusions
The OPENXRDbenchmark provides a controlled and extensible framework for evaluating how large language models (LLMs) and multimodal LLMs integrate external scientific knowledge. Through token-matched experiments and token-budget analyses, the study isolates the effect of content quality from text quantity and reveals that expert-reviewed supporting materials consistently enhance model accuracy, even when passage length is identical. Small and mid-sized models both exhibit clear accuracy improvements when augmented with expert-reviewed materials, with the relative gain being most pronounced for small models, whereas large models show minimal or slightly negative changes due to knowledge saturation. These findings demonstrate that as model capacity increases, the relative advantage of external context diminishes, even though overall accuracy continues to scale with size.
Beyond these controlled experiments, OPENXRD demonstrates strong scalability and generalizability across 74 diverse model architectures, including general-purpose systems (GPT, Claude), reasoning-optimized models (O-series), code-specialized variants (QWEN-coder), domain-adapted frameworks (LLaMAT, dZiner), and vision-language models (LLaVA). This diversity confirms that OPENXRD is model-agnostic, functioning as a diagnostic layer for both closed-book and retrieval-augmented reasoning systems. For instance, the code-specialized model QWEN3-coder-plus improved dramatically, from 22.99% in the closed-book setting (Table 2) to 88.94% in the open-book setting (Table 4) with expert-reviewed materials, demonstrating that OPENXRD’s insights extend beyond language-only systems and generalize effectively across diverse architectural paradigms. Moreover, because OPENXRD isolates inference-time reasoning while remaining compatible with dynamic retrieval pipelines, it can be directly extended into a deployable RAG-evaluation suite that separates retrieval quality from assimilation capability.
We therefore do not report a single “standard RAG baseline” in this work, because end-to-end RAG accuracy depends strongly on retriever choice, corpus construction, chunking, and ranking, which would introduce additional confounds beyond the generator-side assimilation question studied here. Instead, OPENXRD provides the oracle-context upper bound and a controlled harness that can later be paired with retrieved chunks to decompose end-to-end RAG performance into retrieval quality versus generator assimilation.
Finally, the framework’s practical deployment insights show how these findings can guide real-world applications. OPENXRD identifies mid-sized models (7B–70B parameters) augmented with expert-reviewed passages as an optimal balance between cost and performance, approaching or even matching large-model accuracy while requiring significantly fewer computational resources. This provides a cost-effective deployment pathway for organizations with limited budgets. The framework also enables the design of adaptive token-budget strategies that scale with task complexity, along with pre-processing approaches that reduce token usage while preserving essential information. Such model-selection and resource-allocation guidelines directly address the economic and computational realities of deploying AI systems at scale.
In conclusion, OPENXRD is not an artificial or static experiment, but a reproducible, extensible, and practically informative framework. It quantifies how model capacity, resource budgets, and content quality interact under real-world constraints, provides scalable and legally compliant data generation methods, and delivers empirical evidence to support efficient design of open-book and retrieval-augmented reasoning systems. The framework’s diagnostic capabilities empower organizations to make informed decisions about when to scale model size versus when to enhance smaller models with curated external knowledge, thereby advancing both the scientific understanding and the practical deployment of domain-specialized language models.